### v2

# New simple and accurate approximations for quintessence

Abstract
We derive new simple approximations for quintessence solutions to Einsteins’s field equations in spatially homogeneous, isotropic and flat spacetimes. The approximations are typically an order of magnitude more accurate than anything available in the literature, which from an observational perspective makes numerical calculations superfluous. For example, our tracking quintessence approximation yields a $\sim 0.1\%$ maximum relative error of $H(z)/H_{0}$ for the observationally viable inverse power law scalar field potentials, and similarly for viable thawing quintessence models using two slow-roll parameters. We situate the new results in a broader dark energy context, where the present parametrized deviations from $\Lambda$CDM offer simple alternatives to the CPL parametrization, useful for observations at increasingly large redshifts.
1 Introduction
A growing number of increasingly precise cosmological observations of, e.g., type Ia supernovae [1, 2, 3], the cosmic microwave background (CMB) [4], various measurements of the Hubble parameter [5, 6, 7], the universe large-scale structure (LSS) [8, 9], and baryon acoustic oscillations (BAO) [10, 11, 12, 13], suggests that the universe is presently accelerating. In addition cosmological observations indicate that most of the matter in the universe is invisible ‘dark matter’ [14, 15, 16, 17]. The standard interpretation of the observations rely on that General Relativity is an accurate description of gravity on cosmological scales and that the observable Universe is almost spatially homogeneous, isotropic and flat on sufficiently large spatial scales. This leads to that the acceleration of the universe requires a ‘dark energy’ (DE) density component, $\rho_{\mathrm{DE}}$, with sufficiently large negative pressure, while the dark matter seems to be best described by a dominant cold dark matter (CDM) energy density component, $\rho_{\mathrm{m}}$, without pressure. The most simple and remarkably successful model that is compatible with these assumptions is the $\Lambda$CDM model, for which the energy density $\rho_{\mathrm{DE}}$ and pressure $p_{\mathrm{DE}}$ are given by $\rho_{\mathrm{DE}}=-p_{\mathrm{DE}}=\Lambda>0$, where $\Lambda$ is the cosmological constant.
There are, however, theoretical problems, e.g., the nature of dark matter [18, 19], the observed small value of $\Lambda$, the coincidence problem, see e.g. [20] and references therein. Recent tensions between some data sets [21, 22, 23] and that large scale structures seem to form surprisingly early in the universe [24] have further aggravated the situation, which has contributed as motivation for numerous dynamical DE models for which $\rho_{\mathrm{DE}}$ is evolving. Many of these models are purely cosmographically motivated, without any tie to any theory, notably the Chevallier-Polarski-Linder (CPL) parametrization [25, 26] of the DE equation of state parameter $w_{\mathrm{DE}}=p_{\mathrm{DE}}/\rho_{\mathrm{DE}}$ and a plethora of variations of it, see e.g. [27] and references therein.
Some DE models, however, endeavour to have some theoretical motivation, where one of the simplest attempts is quintessence, i.e. a canonical scalar field, $\phi$, that is minimally coupled to gravity, see e.g. [28] and [29] for references. This is the main topic of the present paper, which deals with a non-interacting scalar field and matter in a spatially homogeneous, isotropic and flat spacetime, although note that the ideas in this paper are applicable to many other models and alternative theories. More specifically, we will derive simple and accurate approximations for thawing and tracking quintessence solutions, [^1] According to Tsujikawa (2013) [28] there are three types of quintessence in the literature: thawing, tracking and scaling freezing quintessence, see [30] for subtleties. In this paper we consider tracking quintessence for asymptotically inverse power law scalar field potentials $V\propto\phi^{-p}$, $p>0$, when $\phi\rightarrow 0$. Since scaling freezing quintessence does not yield past continuous deformations of $\Lambda$CDM we will only briefly touch on this topic in appendix B. which yield continuous deformations of $\Lambda$CDM cosmology. Notably, the approximations are typically simpler and an order of magnitude more accurate than anything available in the literature. For example, our tracking quintessence approximation yields a $\sim 0.1\%$ maximum relative error of $H(z)/H_{0}$ for the observationally viable inverse power law scalar field potentials, and similarly for viable thawing quintessence models using two slow-roll parameters. We note that from an observational perspective the remarkable accuracy of our approximations makes numerical calculations superfluous when assessing observational constraints; moreover the approximations yield an analytic overview of the observational viability of broader ranges of models than those that are typically numerically investigated. For example, it would be interesting to replace the recent numerical observational constraint analysis by Shlivko and Steinhardt [31] with our thawing quintessence approximations for a broader set of models, although we note that our new parametrized DE deviations from $\Lambda$CDM offer simple alternatives to the CPL parametrization, useful for observations at increasingly large redshifts.
The outline of the paper is as follows. Motivated by our new quintessence approximations, the next section provides an associated general DE context. In Section 3 we derive the expressions for the parameters appearing in the previous DE section for thawing and tracking quintessence. The new thawing and tracking quintessence approximations are compared with numerical calculations and previous results in the literature in section 4. The main part of the paper ends with section 5, which includes suggestions for future applications of the present approach and results. Appendix A reviews quintessence approximations found in the literature, while appendix B deals with scaling freezing quintessence.
2 Dark energy considerations
2.1 Preliminaries
Spatially homogeneous, isotropic and flat spacetimes have a line element that can be written as [^2] We use units such that $c=1$ and $18\pi G=$ 1, where $c$ is the speed of light and $G$ is the gravitational constant.
$\begin{split}ds^{2}&=-dt^{2}+a^{2}(t)\delta_{ij}dx^{i}dx^{j}\\ &=(1+z)^{-2}\left[H^{-2}(z)dz^{2}+a_{0}^{2}\delta_{ij}dx^{i}dx^{j}\right],\end{split}$ (1)
where $a(t)$ is the cosmological scale factor, $a_{0}=a(t_{0})$ (throughout, a subscript 0 refers to quantities at the present time $t=t_{0}$), where $a_{0}$ is usually set to one by an appropriate scaling of the spatial coordinates;
$H=\frac{\dot{a}}{a}$ (2)
denotes the Hubble parameter, where the dot refers to differentiation with respect to the cosmic time $t$, while $z=(a_{0}/a)-1$ is the redshift.
Since the spacetime is described by the Hubble parameter $H(z)$ it follows that cosmological observables and their interpretation depend directly on $H(z)$, its derivatives and integrals. This is illustrated by, e.g., the deceleration parameter $q=-1-\dot{H}/H^{2}=-1+(1+z)\frac{dH}{dz}/H$, and many distance measures, e.g., the Hubble distance $D_{H}(z)$, the proper motion or coordinate distance $D_{M}(z)$, the luminosity distance $D_{L}(z)$, the angular diameter distance $D_{A}(z)$, which are given by (see, e.g. [27])
$D_{H}$ $=\frac{1}{H(z)},$ (3a)
[…]
To find interesting alternatives to $\Lambda$CDM some authors simply choose some $H(z)$, e.g. $H^{2}$ being a polynomial in $1+z$ as in [27], or obtain $H(z)$ by assuming a specific form for one of the distance measures, e.g. $D_{L}(z)$, which yield $H(z)$ via $d[D_{L}(z)/(1+z)]/dz=1/H(z)$, see e.g. [27]. More commonly, however, is to make a choice of the DE equation of state parameter
$w_{\mathrm{DE}}=\frac{p_{\mathrm{DE}}}{\rho_{\mathrm{DE}}}$ (4)
that is different than the $\Lambda$CDM value $w_{\mathrm{DE}}=-1$ (see e.g. [23] and references therein for a plethora of examples of $w_{\mathrm{DE}}(z)$ and $w_{\mathrm{DE}}(a)$). Note, however, that $w_{\mathrm{DE}}$ is not a physical observable. Since cosmological observations depend on $H(z)$, it is necessary to compute $H(z)$ from $w_{\mathrm{DE}}(z)$, which requires the Einstein field equations and the matter conservation equations.
Although the distance measures are naturally expressed in terms of the redshift $z$, computations are usually simpler using the $e$-fold time $N=\ln(a/a_{0})$, where it is subsequently trivial to transform results to the redshift $z$ by using the relation
$1+z=\exp(-N).$ (5)
For simplicity we consider only non-interacting pressure free matter and a dark energy component. [^3] Although the gravitational influence of radiation is negligible when quintessence is important for the evolution of the Universe, we will comment on how to add it to our approximations in section 5. The matter conservation equations for the matter energy density, $\rho_{\mathrm{m}}>0$ and the DE energy density, $\rho_{\mathrm{DE}}>0$, are given by
$\rho^{\prime}_{\mathrm{m}}$ $=-3\rho_{\mathrm{m}}$$\Rightarrow$ $\quad\rho_{\mathrm{m}}$ $=\rho_{\mathrm{m}0}\exp(-3N),$ (6a)
$\rho^{\prime}_{\mathrm{DE}}$ $=-3(1+w_{\mathrm{DE}})\rho_{\mathrm{DE}}$$\Rightarrow$ $\quad\rho_{\mathrm{DE}}$ $=\rho_{\mathrm{DE}0}\exp(-3N)f(N),$ (6b)
where a denotes the derivative with respect to $N$ and where $f(N)$ is defined as
[…]
2.2 Relationships between $\Omega_{\mathrm{DE}}$, $\Omega_{\mathrm{m}}$, $w_{\mathrm{DE}}$ and $H$
Next we introduce the dimensionless and bounded Hubble normalized variables $\Omega_{\mathrm{DE}}$ and $\Omega_{\mathrm{m}}$:
$\Omega_{\mathrm{DE}}=\frac{\rho_{\mathrm{DE}}}{3H^{2}}=\frac{\rho_{\mathrm{DE}}}{\rho},\qquad\Omega_{\mathrm{m}}=\frac{\rho_{\mathrm{m}}}{3H^{2}}=\frac{\rho_{\mathrm{m}}}{\rho},$ (8)
where we have used the Friedmann equation (i.e. the Gauss constraint)
$3H^{2}=\rho=\rho_{\mathrm{m}}+\rho_{\mathrm{DE}},$ (9)
which yields [^4] To improve our approximations we will use $\Omega_{\mathrm{DE}}(N)$, but subsequently the relation $\Omega_{\mathrm{DE}0}=1-\Omega_{\mathrm{m}0}=1-\rho_{\mathrm{m}0}/3H_{0}^{2}$ can be used to bring the approximations closer to present day data.
$\Omega_{\mathrm{m}}+\Omega_{\mathrm{DE}}=1\quad\Rightarrow\quad\Omega_{\mathrm{m}0}+\Omega_{\mathrm{DE}0}=1.$ (10)
To display the relationship between these quantities and $H$ it is convenient to introduce the present time normalized Hubble parameter
$E=\frac{H}{H_{0}}.$ (11)
Equations (9), (6) and (8) result in the following DE expression for $E^{2}=E_{\mathrm{DE}}^{2}$:
$E_{\mathrm{DE}}^{2}=\Omega_{\mathrm{m}0}\exp(-3N)+\Omega_{\mathrm{DE}0}\exp(-3N)f(N).$ (12)
Using $\rho_{\mathrm{DE}}/\rho_{\mathrm{m}}=\Omega_{\mathrm{DE}}/\Omega_{\mathrm{m}}=\Omega_{\mathrm{DE}}/(1-\Omega_{\mathrm{DE}})$ and (6), (10) yield
$\Omega_{\mathrm{DE}}=\frac{\Omega_{\mathrm{DE}0}}{[\Omega_{\mathrm{m}0}/f(N)]+\Omega_{\mathrm{DE}0}},\qquad\Omega_{\mathrm{m}}=\frac{\Omega_{\mathrm{m}0}}{\Omega_{\mathrm{m}0}+\Omega_{\mathrm{DE}0}f(N)}.$ (13)
Since the above expressions involve $f(N)$, the integral (7) needs to be computed from $w_{\mathrm{DE}}(N)$.
If one instead of $w_{\mathrm{DE}}(N)$ has $\Omega_{\mathrm{DE}}(N)$, then one can obtain $w_{\mathrm{DE}}(N)$ and $E_{\mathrm{DE}}^{2}(N)$ since
$w_{\mathrm{DE}}$ $=\frac{1}{3}\left[\ln\left(\frac{\Omega_{\mathrm{m}}}{\Omega_{\mathrm{DE}}}\right)\right]^{\prime}=\frac{1}{3}\left[\ln\left(\frac{1-\Omega_{\mathrm{DE}}}{\Omega_{\mathrm{DE}}}\right)\right]^{\prime},$ (14a)
$E_{\mathrm{DE}}^{2}$ $=\exp(-3N)\left(\frac{\Omega_{\mathrm{m}0}}{\Omega_{\mathrm{m}}}\right)=\exp(-3N)\left(\frac{1-\Omega_{\mathrm{DE}0}}{1-\Omega_{\mathrm{DE}}}\right),$ (14b)
follows from (6) and (8).
We finally note that the deceleration parameter $q$ is defined as
$q=-\frac{\ddot{a}}{aH^{2}}=-1-\frac{H^{\prime}}{H},$ (15)
and is related to the total equation of state parameter $w_{\mathrm{tot}}$ according to
$1+w_{\mathrm{tot}}=\frac{2}{3}(1+q),$ (16)
where
$w_{\mathrm{tot}}=\frac{p}{\rho}=\frac{p_{\mathrm{DE}}}{\rho}=w_{\mathrm{DE}}\left(\frac{\rho_{\mathrm{DE}}}{\rho}\right)=w_{\mathrm{DE}}\Omega_{\mathrm{DE}}.$ (17)
To derive the above relations, we have used the Friedmann equation $\rho=3H^{2}$ to obtain the last equality in (17), and also (15) and (6) to establish $\rho^{\prime}=-3(1+w_{\mathrm{tot}})\rho$, which also follows from the Einstein field equations.
In the following sections we will discuss quintessence for which $w_{\mathrm{DE}}=w_{\phi}$. As shown in section 3, see also e.g. [30] and [32], thawing quintessence yields $\lim_{N\rightarrow-\infty}w_{\phi}=-1$ while tracking quintessence leads to $\lim_{N\rightarrow-\infty}w_{\phi}=w_{\infty}=\mathrm{constant}$, $-1<w_{\infty}<0$. As a consequence thawing and tracking quintessence are asymptotically toward the past described by the $\Lambda$CDM and $w$CDM models, respectively. Let us therefore recall the properties of these models.
2.3 The $\Lambda$CDM and $w$CDM models
The $\Lambda$CDM and the $w$CDM models are characterized by $w_{\mathrm{DE}}=-1$ and $w_{\mathrm{DE}}=\mathrm{constant}=w$, respectively. This leads to that (7) results in
$f=e^{3N}\quad\text{and}\quad f=e^{-3wN},$ (18)
respectively, where (13) and (12) then yield
$\Omega_{\Lambda}$ $=\frac{\Omega_{\Lambda 0}}{\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{\Lambda 0}},$ $\quad\Omega_{w}$ $=\frac{\Omega_{w0}}{\Omega_{\mathrm{m}0}e^{3wN}+\Omega_{w0}},$ (19a)
$E^{2}_{\Lambda}$ $=\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{\Lambda 0},$ $\quad E^{2}_{w}$ $=\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{w0}e^{-3(1+w)N},$ (19b)
where we use the subscripts Λ and w to characterize the $\Lambda$CDM and $w$CDM expressions for $\Omega_{\mathrm{DE}}$ and $E_{\mathrm{DE}}^{2}$, respectively.
2.4 Past DE series expansions
Thawing and tracking quintessence, as illustrated in a dynamical systems setting in [30] and [32], are associated with attractor solutions that originate from fixed points on the matter dominant boundary where $\Omega_{\phi}=0$ and $\Omega_{\mathrm{m}}=1$. Treating both cases simultaneously as dark energy models, i.e., by letting $w_{\mathrm{DE}}\rightarrow w_{\infty}\in[-1,0)$ and $\Omega_{\mathrm{DE}}\rightarrow 0$ when $N\rightarrow-\infty$, leads to that equation (14a), which can be written as $\Omega_{\mathrm{DE}}^{\prime}=-3w_{\mathrm{DE}}\Omega_{\mathrm{DE}}(1-\Omega_{\mathrm{DE}})$, results in the approximation $\Omega_{\mathrm{DE}}^{\prime}=-3w_{\infty}\Omega_{\mathrm{DE}}$ in this limit and hence
$\Omega_{\mathrm{DE}}\approx Ce^{-3w_{\infty}N},$ (20)
which is indeed the case for the thawing ($w_{\infty}=-1$) and tracking ($w_{\infty}\in(-1,0)$) quintessence models. To improve the approximation we now perform series expansions in
$T=T_{0}e^{-3w_{\infty}N},$ (21)
where we for brevity only consider $w_{\mathrm{DE}}$ and $\Omega_{\mathrm{DE}}$ (note that $T\rightarrow 0$ when $N\rightarrow-\infty$). If one starts with a series expansion in $w_{\mathrm{DE}}$ and uses (7) and (13) one obtains an expression for $\Omega_{\mathrm{DE}}$ that subsequently can be series expanded in $T$; on the other hand, if one starts with a series expansion in $\Omega_{\mathrm{DE}}$ and inserts this into (14a) one obtains an expression for $w_{\mathrm{DE}}$ that one then can series expand in $T$. These two procedures result in the same series expansions, which can be written as [^5] The given expressions can be generalized by writing $w_{\mathrm{DE}}=w_{\infty}\left[1+\alpha T+\sigma T^{2}+\dots\right]$, which subsequently can be used to compute the series expansions for $\Omega_{\mathrm{DE}}$ and, e.g., $E_{\mathrm{DE}}^{2}$. The reason we have written $\alpha$ as $1-\gamma$ and $\sigma$ as $(\gamma-1)\beta$ is that this simplifies and is compatible with our thawing and tracking quintessence approximations, where these use the $[1/1]$ Padé expression for $\Omega_{\mathrm{DE}}$ in order to connect our approximations to the present time.
$w_{\mathrm{DE}}$ $=w_{\infty}\left[1-(\gamma-1)T+(\gamma-1)\beta T^{2}+\dots\right]$ (22a)
$\Omega_{\mathrm{DE}}$ $=T\left(1-\gamma T+\left[1+\frac{1}{2}(\gamma-1)\left(3+\gamma+\beta\right)\right]T^{2}+\dots\right).$ (22b)
2.5 DE Padé approximants
For brevity we will restrict our considerations by following the most common approach in the literature, which is to use $w_{\mathrm{DE}}$ as the starting point. The series expansions provide approximations when $T$ is small, i.e. close to the matter dominated regime. To obtain approximations that are accurate even up to the present time we will use Padé approximants for the series expansion of $w_{\mathrm{DE}}$ in $T$. Our first and second order expansions in $T$ for $w_{\mathrm{DE}}$ yield the following Padé approximants for $w_{\mathrm{DE}}$:
$w_{\mathrm{DE}}$ $\approx[0/1]_{w_{\mathrm{DE}}}(T)=\frac{w_{\infty}}{1+(\gamma-1)T},$ (23a)
$w_{\mathrm{DE}}$ $\approx[1/1]_{w_{\mathrm{DE}}}(T)=w_{\infty}\left(1-\frac{(\gamma-1)T}{1+\beta T}\right).$ (23b)
To connect with the present time we can set $N=0$ and solve for $T=T_{0}$ in (23) in terms of the left hand side $w_{\mathrm{DE}0}$. However, since $w_{\mathrm{DE}0}$ is not an observable we will instead relate $T_{0}$ to $\Omega_{\mathrm{DE}0}$ and thereby obtain a direct connection to the observables $H_{0}$ and $\rho_{\mathrm{m}0}$ since $1-\Omega_{\mathrm{DE}0}=\Omega_{\mathrm{m}0}=\rho_{\mathrm{m}0}/3H_{0}^{2}$. To obtain accurate approximations at the present time for $\Omega_{\mathrm{DE}}$ we consider the Padé approximant (use the expression (22b) up to second order in $T$)
$\Omega_{\mathrm{DE}}\approx[1/1]_{\Omega_{\mathrm{DE}}}(T)=\frac{T}{1+\gamma T},$ (24)
from which it follows that $\Omega_{\mathrm{DE}0}=T_{0}/[1+\gamma T_{0}]$, which yields
$T_{0}=\frac{\Omega_{\mathrm{DE}0}}{1-\gamma\Omega_{\mathrm{DE}0}}.$ (25)
This is then inserted into (23a), which when used in (7) to obtain $f(N)$ leads to the same expression as (24) with (25) inserted. To obtain a better approximation for $\Omega_{\mathrm{DE}}(N)$ one can then go on to Padé approximants also involving $\beta$ in the expansion. [^6] These are given by $[1/2]_{\Omega_{\mathrm{DE}}}=\frac{T}{1+\gamma T-\frac{1}{2}(\gamma-1)\left[1+(\beta-\gamma)\right]T^{2}}$ and $[2/1]_{\Omega_{\mathrm{DE}}}=\frac{\gamma T+\frac{1}{2}(\gamma-1)\left[1+(\beta-\gamma)\right]T^{2}}{\gamma+\left[1+\frac{1}{2}(\gamma-1)\left(3+\beta+\gamma\right)\right]T}$. To express $T_{0}$ in terms of $\Omega_{\mathrm{DE}0}$ obtained in this manner requires solving quadratic equations in $T_{0}$, which leads to rather messy expressions. The approximation (24) turns out to be a fairly good approximation for quintessence. For this reason we will use the simple result in (25) for $T_{0}$ in (23b). This leads to that the expressions for $w_{\mathrm{DE}}$ in (23) can be written as follows: [^7] Note that one can use the series expansion in $T$ for $\Omega_{\mathrm{DE}}$ to obtain the following expression for $E_{\mathrm{DE}}^{2}$: $E_{\mathrm{DE}}^{2}\approx\Omega_{\mathrm{m}0}\exp(-3N)[1+T-(\gamma-1)T^{2}(1+{\textstyle{1\over 2}}(\gamma-1+\beta)T)+\dots]$ and use (25) for $T_{0}$. This expression can subsequently be used to obtain various Padé approximants to obtain good quintessence approximations for $E_{\mathrm{DE}}=E_{\phi}$ and then use $E_{\phi}(N)$ to compute, e.g., $w_{\mathrm{DE}}=w_{\phi}$, if one is so inclined.
$w_{\mathrm{DE}}$ $\approx[0/1]_{w_{\mathrm{DE}}}(T)=w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{(\gamma-1)\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})e^{3w_{\infty}N}}\right),$ (26a)
$w_{\mathrm{DE}}$ $\approx[1/1]_{w_{\mathrm{DE}}}(T)=w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{\beta\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})e^{3w_{\infty}N}}\right),$ (26b)
which result in the following present time values
$w_{\mathrm{DE}0}$ $\approx w_{\infty}\left(\frac{1-\gamma\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{m}0}}\right),$ (27a)
$w_{\mathrm{DE}0}$ $\approx w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{1+(\beta-\gamma)\Omega_{\mathrm{DE}0}}\right).$ (27b)
Note that setting $\beta=\gamma-1$ in $[1/1]_{w_{\mathrm{DE}}}(T)$ yields $[0/1]_{w_{\mathrm{DE}}}(T)$ and that setting $\gamma=1$ in both $[0/1]_{w_{\mathrm{DE}}}(T)$ and $[1/1]_{w_{\mathrm{DE}}}(T)$ results in $w_{\mathrm{DE}}=w_{\infty}$, i.e., $\gamma=1$ leads to the $w$CDM model when $w_{\infty}\in(-1,0)$ and the $\Lambda$CDM model when $w_{\infty}=-1$.
Inserting the expressions for $[0/1]_{w_{\mathrm{DE}}}(T)$ and $[1/1]_{w_{\mathrm{DE}}}(T)$ in (23) and (26) into (7) yields
$f$ $=\frac{T}{T_{0}}\left(\frac{1+(\gamma-1)T_{0}}{1+(\gamma-1)T}\right)=\frac{\Omega_{\mathrm{m}0}e^{-3w_{\infty}N}}{1-\gamma\Omega_{\mathrm{DE}0}+(\gamma-1)\Omega_{\mathrm{DE}0}e^{-3w_{\infty}N}},$ (28a)
$f$ $=\frac{T}{T_{0}}\left(\frac{1+\beta T_{0}}{1+\beta T}\right)^{\frac{\gamma-1}{\beta}}=e^{-3w_{\infty}N}\left(\frac{1+(\beta-\gamma)\Omega_{\mathrm{DE}0}}{1-\gamma\Omega_{\mathrm{DE}0}+\beta\Omega_{\mathrm{DE}0}e^{-3w_{\infty}N}}\right)^{\frac{\gamma-1}{\beta}},$ (28b)
where the first $f$ comes from $[0/1]_{w_{\mathrm{DE}}}(T)$ and the second from $[1/1]_{w_{\mathrm{DE}}}(T)$.
It follows from (13) and (12) that the $[1/1]_{w_{\mathrm{DE}}}(T)$ Padé approximant results in
$\Omega_{\mathrm{DE}}$ $\approx\frac{\Omega_{\mathrm{DE},0}}{\Omega_{\mathrm{m}0}e^{3w_{\infty}N}\left(\frac{1+\beta T}{1+\beta T_{0}}\right)^{\frac{\gamma-1}{\beta}}+\Omega_{\mathrm{DE},0}},$ (29a)
$E_{\mathrm{DE}}^{2}$ $\approx\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{\mathrm{DE}0}e^{-3(1+w_{\infty})N}\left(\frac{1+\beta T_{0}}{1+\beta T}\right)^{\frac{\gamma-1}{\beta}}$ (29b)
where we recall that $T=T_{0}\exp(-3w_{\infty}N)$, $T_{0}=\Omega_{\mathrm{DE}0}/(1-\gamma\Omega_{\mathrm{DE}0})$, while the result for the $[0/1]_{w_{\mathrm{DE}}}(T)$ Padé is obtained from the above expressions by setting $\beta=\gamma-1$, which yields [^8] Note that by multiplying (30a) with $\gamma$ it is easy to see that the transformation $\gamma\Omega_{\mathrm{DE}}\rightarrow\Omega_{w}$ and $w_{\infty}\rightarrow w$ yields the expression in (19a), which specializes to $\Lambda$CDM when $w_{\infty}=-1$. In particular this leads to $\gamma\Omega_{\mathrm{DE}0}=\Omega_{w0}$, which illustrates the danger of interpreting and transferring data from a $w$CDM or $\Lambda$CDM context to a different DE context.
$\Omega_{\mathrm{DE}}$ $\approx\frac{\Omega_{\mathrm{DE},0}}{\gamma\Omega_{\mathrm{DE},0}+(1-\gamma\Omega_{\mathrm{DE},0})e^{3w_{\infty}N}},$ (30a)
$E_{\mathrm{DE}}^{2}$ $\approx\Omega_{\mathrm{m}0}e^{-3N}\left(1+\frac{\Omega_{\mathrm{DE}0}}{(\gamma-1)\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})e^{3w_{\infty}N}}\right).$ (30b)
It is trivial to express the above relations in the redshift $z$, which is done explicitly in section 4 where we investigate the accuracy of thawing and tracking quintessence approximations. To do so, we need $\gamma$ and $\beta$, which we derive in the next section, but before doing this let us briefly consider the CPL parametrization.
2.6 The Chevallier-Polarski-Linder parametrization
The so-called CPL parameters [25, 26] are often used in cosmology [^9] For a recent discussion of the CPL parametrization, see Wolf & Ferreira [33]. and parametrize $w_{\mathrm{DE}}$ according to
$w_{\mathrm{DE}}=w_{\mathrm{CPL}}=w_{0}+w_{a}\left(\frac{z}{1+z}\right)=w_{\infty}-w_{a}e^{N},$ (31)
where
$w_{\infty}=w_{0}+w_{a}.$ (32)
By using that $w_{a}=-w_{\mathrm{DE}}^{\prime}|_{N=0}$ we can easily compute the parameter $w_{a}$ from (26):
$w_{a}$ $=\frac{-3(\gamma-1)w_{\infty}^{2}\Omega_{\mathrm{DE}0}(1-\gamma\Omega_{\mathrm{DE}0})}{\Omega_{\mathrm{m}0}^{2}},$ (33a)
$w_{a}$ $=\frac{-3(\gamma-1)w_{\infty}^{2}\Omega_{\mathrm{DE}0}(1-\gamma\Omega_{\mathrm{DE}0})}{(1+(\beta-\gamma)\Omega_{\mathrm{DE}0})^{2}}.$ (33b)
Inserting (31) into (7) yields
$f=e^{-3w_{\infty}N}\cdot e^{-3w_{a}(1-e^{N})}=(1+z)^{3w_{\infty}}\cdot e^{-3w_{a}\left(\frac{z}{1+z}\right)},$ (34)
which results in that (13) and (12) yield
$\Omega_{\mathrm{CPL}}$ $=\frac{\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{DE}0}+\Omega_{\mathrm{m}0}e^{3w_{\infty}N}\cdot e^{3w_{a}(1-e^{N})}},$ (35a)
$E_{\mathrm{CPL}}^{2}$ $=\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{\mathrm{DE}0}e^{-3(1+w_{\infty})N}\cdot e^{-3w_{a}(1-e^{N})},$ (35b)
and thereby
$\Omega_{\mathrm{CPL}}$ $=\frac{\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{DE}0}+\Omega_{\mathrm{m}0}(1+z)^{-3w_{\infty}}\cdot e^{3w_{a}\left(\frac{z}{1+z}\right)}},$ (36a)
$E_{\mathrm{CPL}}^{2}$ $=\Omega_{\mathrm{m}0}(1+z)^{3}+\Omega_{\mathrm{DE}0}(1+z)^{3(1+w_{\infty})}\cdot e^{-3w_{a}\left(\frac{z}{1+z}\right)}.$ (36b)
One can regard (31) as a series expansion in $e^{N}$, but this is inconsistent with the past matter dominant requirement $\Omega_{\mathrm{DE}}\rightarrow 0$ and that one of the CPL parameters can be interpreted as $w_{\infty}$, since this naturally leads to an expansion in $e^{-3w_{\infty}N}$, which is particularly pertinent as observations probe increasingly large redshift. CPL thus mixes this asymptotically motivated function (arising from the $w_{\infty}$ part in (31)) in $\Omega_{\mathrm{CPL}}$ and $E_{\mathrm{CPL}}^{2}$ with another function that comes from the $w_{a}$ part in (31). Due to this mixture of functions, especially the exponential part involving $\frac{z}{1+z}$, the CPL parametrization arguably does not lead to simpler expressions for observables than our $[0/1]_{w_{\mathrm{DE}}}(T)$ and $[1/1]_{w_{\mathrm{DE}}}(T)$ based results, and it certainly is not a good approximation for thawing and tracking quintessence.
3 Deriving the parameters $\gamma$ and $\beta$ for quintessence
3.1 Thawing quintessence
To obtain a dynamical system that is useful for deriving expressions for $\gamma$ and $\beta$ for thawing quintessence [^10] The concepts of thawing and freezing were defined by Caldwell and Linder (2005) [34] as follows: thawing is characterized by $w_{\varphi}\approx-1$ where $w_{\varphi}$ subsequently grows, i.e. $w_{\varphi}^{\prime}>0$, while $w_{\varphi}>-1$ and $w_{\varphi}^{\prime}<0$ holds for freezing. we introduce the following dimensionless bounded variable [^11] The variable $\Sigma_{\phi}$ was first introduced by Coley et al. (1997) [35] and Copeland et al. (1998) [36] whose $x$ is $\Sigma_{\varphi}$. Since then, $\Sigma_{\phi}$ (or $\phi^{\prime}$) is commonly used to describe scalar fields in cosmology, see, e.g., Urena-Lopez (2012) [37], equation (2.3), Tsujikawa (2013) [28], equation (16) and Alho and Uggla (2015) [38], equation (8). The reason for using the notation $\Sigma$ for the kernel is because $\Sigma_{\phi}$ plays a similar role as Hubble-normalized shear, which is typically denoted with the kernel $\Sigma$, see e.g. [39].
$\Sigma_{\phi}=\frac{\phi^{\prime}}{\sqrt{6}},$ (37)
and use $({\phi},\Sigma_{\phi},\Omega_{\phi})$ as the state vector together with the $e$-fold time $N$ [^12] The variables $u$ and $v$ in [30] yield slower convergence. To obtain a dynamical system for $({\phi},\Sigma_{\phi},\Omega_{\phi})$ we use that (14a) is equivalent to $\Omega_{\mathrm{DE}}^{\prime}=-3w_{\mathrm{DE}}\Omega_{\mathrm{DE}}(1-\Omega_{\mathrm{DE}})$, $\Omega_{\mathrm{DE}}=\Omega_{\phi}$, and that $w_{\mathrm{DE}}=w_{\phi}$ is given by
$w_{\phi}=-1+\frac{2\Sigma_{\phi}^{2}}{\Omega_{\phi}},$ (38)
as follows from the definitions. In addition, the (non-linear) Klein-Gordon equation $\ddot{\varphi}=-3H\dot{\varphi}-V_{,\varphi}$, expressed in $e$-fold time and $\Sigma_{\phi}$, is needed. This leads to the dynamical system
${\phi}^{\prime}$ $=\sqrt{6}\Sigma_{\phi},$ (39a)
$\Sigma_{\phi}^{\prime}$ $=-\frac{3}{2}(1+\Omega_{\phi}-2\Sigma_{\phi}^{2})\Sigma_{\phi}+\sqrt{\frac{3}{2}}\,\lambda({\phi})(\Omega_{\phi}-\Sigma_{\phi}^{2}),$ (39b)
$\Omega^{\prime}_{\phi}$ $=3(\Omega_{\phi}-2\Sigma_{\phi}^{2})(1-\Omega_{\phi}),$ (39c)
where we have introduced
$\lambda(\phi)=-\frac{V_{,\phi}}{V}.$ (40)
Since thawing quintessence begins in the matter dominated regime, $\Omega_{\mathrm{m}}\approx 1$, when $w_{\varphi}\approx-1$ initially, it follows from (38) and from $\Omega_{\mathrm{m}}+\Omega_{\phi}=1$ that thawing quintessence begins when $\Sigma_{\varphi}^{2}\ll\Omega_{\phi}\ll\Omega_{\mathrm{m}}$. We then note that whenever $\lambda(\phi)$ is bounded, which always holds for some domain of $\phi$ for any scalar field potential $V(\phi)$, there is a line of matter dominated ‘Friedmann-Lemaître’ fixed points given by
$\mathrm{FL}^{\phi_{*}}\,:\quad(\phi,\Sigma_{\phi},\Omega_{\phi})=(\phi_{*},0,0)$ (41)
and hence $\Omega_{\mathrm{m}}=1$, parameterized by the constant values $\phi_{*}$. To determine what happens in the vicinity of $\mathrm{FL}^{\phi_{*}}$ we linearize the equations at $(\phi,\Sigma_{\phi},\Omega_{\phi})=(\phi_{*},0,0)$, which leads to
$\phi$ $\approx\phi_{*}+\frac{2\lambda_{*}}{9}Ce^{3N}-\frac{2}{3}Ke^{-\frac{3}{2}N},$ (42a)
$\sqrt{6}\Sigma_{\phi}$ $=\phi^{\prime}\approx\frac{2\lambda_{*}}{3}Ce^{3N}+Ke^{-\frac{3}{2}N}$ (42b)
$\Omega_{\phi}$ $\approx Ce^{3N},$ (42c)
where $\lambda_{*}=\lambda(\phi_{*})$. It follows that thawing quintessence is associated with the unstable manifold of $\mathrm{FL}^{\phi_{*}}$, obtained by setting $K=0$. More precisely, thawing quintessence models consist of the 1-parameter set of unstable manifold orbits (i.e., solution trajectories) of $\mathrm{FL}^{\phi_{*}}$ and the open set of orbits that intermediately shadow these orbits; this open set of orbits consists of the orbits that before the thawing epoch are pushed toward $\mathrm{FL}^{\phi_{*}}$ by the stable manifold orbits of $\mathrm{FL}^{\phi_{*}}$ and subsequently toward the unstable manifold orbits of $\mathrm{FL}^{\phi_{*}}$, which therefore approximate these orbits during their thawing stage of evolution, for details, see [30] and [40].
To find $\gamma$ and $\beta$ we Taylor expand $\lambda(\phi)$ around $\phi_{*}$ and make a series expansion in $T=T_{0}e^{3N}$ for the variables $(\phi,\Sigma_{\phi},\Omega_{\phi})$. The coefficients in the expansion are found by inserting the expansions into the dynamical system (39) and solving for them. Truncation of the series in $T$ gives
$\phi$ $\approx\phi_{*}+\sqrt{(\gamma-1)/3}\left[T\left(1-\frac{1}{2}\sigma T\right)\right],$ (43a)
$\sqrt{6}\Sigma_{\phi}$ $=\phi^{\prime}\approx\sqrt{3(\gamma-1)}\left[T(1-\sigma T)\right],$ (43b)
$\Omega_{\phi}$ $\approx T\left[1-\gamma T+\left(1+(\gamma-1)\left(\frac{3}{2}+\sigma\right)\right)T^{2}\right],$ (43c)
where $\gamma$ and $\sigma$ are related to the slow-roll parameters
$\epsilon(\phi)=\frac{1}{2}\left(\frac{V_{,\phi}}{V}\right)^{2}=\frac{\lambda^{2}}{2},\qquad\eta(\phi)=\frac{V_{,\phi\phi}}{V}=\lambda^{2}-\lambda_{,\phi},$ (44)
computed at $\phi=\phi_{*}$ on $\mathrm{FL}^{\phi_{*}}$, i.e.,
$\epsilon_{*}=\epsilon(\phi_{*}),\qquad\eta_{*}=\eta(\phi_{*}),$ (45)
according to
$\gamma=1+\left(\frac{2}{3}\right)^{3}\epsilon_{*},\qquad\sigma=\frac{4}{5}\left(1+\frac{\eta_{*}}{6}\right);$ (46)
where a comparison between (22b) and (43c) results in
$\beta=\frac{3}{5}-\left(\frac{2}{3}\right)^{3}\epsilon_{*}+\frac{4}{15}\eta_{*}.$ (47)
3.2 Tracking quintessence
Peebles and Ratra (1988) [41, 42] initiated the study of the inverse power law potential $V\propto\phi^{-p}$, $p>0$. They obtained approximate expressions by assuming that $H$ and $\rho$ could be approximated by their values in the matter dominated regime and that the scale factor obeyed a power law dependence in cosmic time $t$. Subsequently Ratra and Quillen (1992) [43] obtained the asymptotic result $w_{\phi}=p_{\phi}/\rho_{\phi}=-2/(2+p)$ (their eq. (5)) for this potential. This work was then followed up by Zlatev et al. (1999) [44] and Steinhardt et al. (1999) [45], [^13] See also Podario and Ratra (2000) [46] and Peebles and Ratra (2003) [20]. who introduced the notation ‘tracker field’ as a form of quintessence, associated with potentials for which $\lambda\rightarrow\infty$ when $\phi\rightarrow 0$ and the condition $\Gamma>1$, see also [28], where
$\Gamma=\frac{V\,V_{,\phi\phi}}{V_{,\phi}^{2}}=1+(\lambda^{-1})_{,\phi}.$ (48)
We are here considering asymptotically inverse power law potentials, characterized by
$\lim_{\phi\rightarrow 0}\phi\lambda=p=\mathrm{constant}>0,\qquad\lim_{\phi\rightarrow 0}\Gamma=1+\frac{1}{p}.$ (49)
Next, we introduce a dynamical system that is suitable for deriving approximations for the evolution in the vicinity of $\phi=0$ and as long the potential and $\lambda$ are monotonically decreasing, i.e., when $\lambda(\phi)>0$ and $\Gamma(\phi)>1$, as follows from (40) and (48). The system of equations are based on the following new variables: [^14] The variables $\tilde{\phi}$ and $v$ are slightly different than the variables $\bar{\varphi}$ and $v$ in [32], which were adapted to a global state space setting, while the present ones are designed for fast convergence for the tracker orbit and the explicit appearance of $\Gamma$ in the equations and results.
$\tilde{\phi}=\lambda^{-2}(\phi),\qquad u=\Sigma_{\phi}\sqrt{\frac{2}{\Omega_{\phi}}}=\frac{\phi^{\prime}}{\sqrt{3\Omega_{\phi}}},\qquad v=\lambda(\phi)\sqrt{\frac{\Omega_{\phi}}{3}},$ (50)
from which it follows, when taken together with (38),
$1+w_{\phi}=u^{2},\qquad\Omega_{\phi}=3v^{2}\tilde{\phi}.$ (51)
These variables lead to the dynamical system
$\tilde{\phi}^{\prime}$ $=6uv\tilde{\phi}(\Gamma(\tilde{\phi})-1),$ (52a)
$u^{\prime}$ $=\frac{3}{2}\left(2-u^{2}\right)(v-u),$ (52b)
$v^{\prime}$ $=\frac{3}{2}\left[(1-u^{2})(1-3v^{2}\tilde{\phi})-2uv(\Gamma(\tilde{\phi})-1)\right]v,$ (52c)
where we assume that $\Gamma(\tilde{\phi})>1$ is regular in $\tilde{\phi}$.
The dynamical system (52) admits $\tilde{\phi}=0$ (and thereby $\phi=0$) as an invariant boundary subset. As shown in [32], in regularizing variables tracking quintessence is associated with a tracker orbit that originates from a tracker fixed point, $\mathrm{T}$, on the $\tilde{\phi}=0$ boundary, which in the present variables is given by
$\mathrm{T}:\qquad(\tilde{\phi},u,v)=(0,u_{\mathrm{T}},v_{\mathrm{T}})=\left(0,1,1\right)\sqrt{\frac{p}{2+p}}\,.$ (53)
Notably $\mathrm{T}$ is a hyperbolic saddle with the ‘tracker orbit’ as its unstable manifold and $\tilde{\phi}=0$ being its 2D stable manifold, which results in that the tracker orbit acts as an ‘attractor orbit’, see [32] for details.
To find the parameters $w_{\infty}$, $\gamma$ and $\beta$ we first note that the tracker fixed point value for $u=u_{\mathrm{T}}$ gives the initial asymptotic tracker value for $w_{\phi}$ since
$w_{\infty}=w_{\phi}|_{\mathrm{T}}=u_{\mathrm{T}}^{2}-1=\frac{p}{2+p}-1=-\frac{2}{2+p},$ (54)
where $w_{\infty}$ describes the single positive eigenvalue, given by $-3w_{\infty}=6/(2+p)$. Next we proceed and make a series expansion in
$T=T_{0}e^{-3w_{\infty}N}$ (55)
for the variables $(\tilde{\phi},u,v)$. Since the governing equations (52) involve $\Gamma(\tilde{\phi})$ we need to use a truncated Taylor series of $\Gamma(\tilde{\phi})$ at $\tilde{\phi}=0$:
$\Gamma(\tilde{\phi})\approx\Gamma^{(0)}+\Gamma^{(1)}\tilde{\phi}+\frac{1}{2}\Gamma^{(2)}\tilde{\phi}^{2}+\dots,\qquad\Gamma^{(n)}:=\left.\frac{d^{n}\Gamma}{d\tilde{\phi}^{n}}\right|_{\tilde{\phi}=0},$ (56)
where $\Gamma^{(0)}=1+p^{-1}$.
To identify $\gamma$ and $\beta$ we solve for the coefficients for the series expanded variables $(\tilde{\phi},u,v)$ by using the equations in (52), which when inserted into $w_{\phi}=u^{2}-1$ yields $w_{\phi}\approx w_{\infty}\left[1-(\gamma-1)T(1-\beta T)\right]$ with
$\gamma$ $=1-w_{\infty}^{-1}(1-w_{\infty}^{2})k,$ (57a)
$\beta$ $=\frac{2w_{\infty}^{2}(3w_{\infty}-1)+k\left(12w_{\infty}^{4}-w_{\infty}^{3}-3w_{\infty}^{2}+2w_{\infty}-1\right)+k^{(2)}}{w_{\infty}(12w_{\infty}^{2}-3w_{\infty}+1)},$ (57b)
$k$ $=\frac{w_{\infty}-\frac{2}{3}\Gamma^{(1)}}{4w_{\infty}^{2}-2w_{\infty}+1},$ (57c)
$k^{(2)}$ $=\frac{w_{\infty}\Gamma^{(2)}}{9(w_{\infty}+1)k},$ (57d)
where we in the expression for $k^{(2)}$ assume that $k\neq 0$. Note that the special class of models with $k=0$, i.e. $\Gamma^{(1)}=\frac{3}{2}w_{\infty}$, needs special treatment and that it include models with $w_{\phi}=\mathrm{constant}=w_{\infty}$, see [32] for details. [^15] Sahni and Starobinsky (2000) [47], equation (121), and Urena-Lopez and Matos (2000) [48], investigated models with the potential $V=V_{0}\sinh^{-p}(\nu\phi)$ and showed that these models admit a special solution (which turns out to be the tracker orbit) for which $w_{\phi}$ is constant, i.e., they yield $w$CDM spacetimes with $-1<w_{\phi}=w_{\infty}=w<0$.
4 Numerical comparisons
In this section we calculate the numerical relative errors of the approximations, i.e.,
$\left|\frac{\Delta F}{F}\right|=\left|\frac{F_{approx}-F_{num}}{F_{num}}\right|,$ (58)
of our approximations for thawing and tracking quintessence, but also for approximations found in the literature, described in appendix A. For brevity we will consider the relative errors of $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)$ and $E(z)$.
Before doing so, let us write our approximations for $w_{\mathrm{DE}}$, $\Omega_{\mathrm{DE}}$, from which $\Omega_{\mathrm{m}}$ follows since $\Omega_{\mathrm{m}}=1-\Omega_{\mathrm{DE}}$, and $E_{\mathrm{DE}}^{2}$ in the redshift $z$. The approximations based on $[0/1]_{w_{\mathrm{DE}}}(T)$ and only involving $\gamma$ are given by (as follows from (26a), (30a) and (30b)):
$w_{\mathrm{DE}}$ $\approx w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{(\gamma-1)\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})(1+z)^{-3w_{\infty}}}\right),$ (59a)
$\Omega_{\mathrm{DE}}$ $\approx\frac{\Omega_{\mathrm{DE},0}}{\gamma\Omega_{\mathrm{DE},0}+(1-\gamma\Omega_{\mathrm{DE},0})(1+z)^{-3w_{\infty}}},$ (59b)
$E_{\mathrm{DE}}^{2}$ $\approx\Omega_{\mathrm{m}0}(1+z)^{3}\left(1+\frac{\Omega_{\mathrm{DE}0}}{(\gamma-1)\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})(1+z)^{-3w_{\infty}}}\right),$ (59c)
while the approximations obtained from $[1/1]_{w_{\mathrm{DE}}}(T)$, involving both $\gamma$ and $\beta$, are as follows:
$w_{\mathrm{DE}}$ $\approx w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{\beta\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})(1+z)^{-3w_{\infty}}}\right),$ (60a)
$\Omega_{\mathrm{DE}}$ $\approx\frac{\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{m}0}(1+z)^{-3w_{\infty}}\left(\frac{1+\beta T}{1+\beta T_{0}}\right)^{\frac{\gamma-1}{\beta}}+\Omega_{\mathrm{DE}0}},$ (60b)
$E_{\mathrm{DE}}^{2}$ $\approx(1+z)^{3}\left[\Omega_{\mathrm{m}0}+\Omega_{\mathrm{DE}0}(1+z)^{3w_{\infty}}\left(\frac{1+\beta T_{0}}{1+\beta T}\right)^{\frac{\gamma-1}{\beta}}\right],$ (60c)
where
$T=T_{0}(1+z)^{3w_{\infty}},\qquad T_{0}=\frac{\Omega_{\mathrm{DE}0}}{1-\gamma\Omega_{\mathrm{DE}0}}.$ (60d)
The parameters $\gamma$ and $\beta$ for thawing quintessence were given in equations (46) and (47), while those for tracking quintessence are found in equations (57), where we recall that in the tracking quintessence case $w_{\infty}=-2/(2+p)$ where $p=\lim_{\phi\rightarrow 0}(\phi\lambda)$. We cannot of course test all possible scalar field models, but we have done so for several potentials and obtained similar results. For brevity we will here only present the results for two scalar field potentials.
Thawing quintessence
For thawing quintessence we will consider a currently popular model, the quintessential $\alpha$-attractor inflation model introduced in [49] and further explored in [50], [51] and [40], referred to as Exp-model I in [50], [51] and as the EC potential model in [40], given by
$V=V_{-}e^{-\nu\left(1+\tanh{\left(\frac{\phi}{\sqrt{6\alpha}}\right)}\right)},$ (61)
which results in
$\begin{split}\lambda&=\frac{\nu}{\sqrt{6\alpha}}\left(1-\tanh^{2}{\left(\frac{\phi}{\sqrt{6\alpha}}\right)}\right),\\ \lambda_{,\phi}&=-\frac{\nu}{3\alpha}\tanh{\left(\frac{\phi}{\sqrt{6\alpha}}\right)}\left(1-\tanh^{2}{\left(\frac{\phi}{\sqrt{6\alpha}}\right)}\right).\end{split}$ (62)
To obtain an explicit model, we set the model parameters $\alpha$ and $\nu$ as follows: $(\alpha,\nu)=(\frac{7}{3},128)$. We considered these models in [40] in the context quintessential inflation, where inflationary considerations determined $V_{-}$, see [50] and [51]. The inclusion of radiation yielded an integral, which resulted in that only one of the thawing solutions in the 1-parameter set characterized by $\phi_{*}$ was compatible with the chosen values of $V_{-}$, $H_{0}$, $\Omega_{\gamma 0}$ and $\Omega_{\mathrm{m}0}$. This solution was found by varying the frozen scalar field value, $\phi_{*}$, at the line of fixed points from which the thawing solutions originated, which resulted in $\phi_{*}=9.594$.
In the present case where we neglect radiation the integral does not exist and then $H_{0}$ and $\Omega_{\phi 0}=0.68$ do not impose any restrictions on $\phi_{*}$ at the line of fixed points $(\phi,\Sigma_{-},\Omega_{\phi})=(\phi_{*},0,0)$ from which the thawing quintessence oribits originate. However, due to that radiation is negligible when quintessence affects the evolution, the present approximations are excellent also when radiation is included as in [40]. For this reason we choose $\phi_{*}=9.594$. Inserting this value into (62) enables the subsequent computation of $\gamma$ and $\beta$ by means of (46) and (47), which results in $(\gamma,\beta)=(1.095,0.789)$, thereby making the approximation numerically explicit. To obtain the relative errors we numerically integrate the dynamical system (39) and the decoupled equation for $E_{\phi}$,
$E_{\phi}^{\prime}=-\frac{3}{2}\left(1+w_{\mathrm{tot}}\right)E_{\phi}=-\frac{3}{2}\left(1+w_{\phi}\Omega_{\phi}\right)E_{\phi}=-\frac{3}{2}\left(1-\Omega_{\phi}+2\Sigma_{\phi}^{2}\right)E_{\phi},$ (63)
where we have used (15), (16) and (38). We then change from $e$-fold time $N$ to the redshift $z$ via $z=\exp(-N)-1$.
The relative errors of $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)$ and $E(z)$ are depicted in figure 1 for our $[0/1]_{w_{\mathrm{DE}}}(T)$ and $[1/1]_{w_{\mathrm{DE}}}(T)$ based approximations, as well as the approximations by Scherrer and Sen [52] and Chiba [53], given by the equations (70) and (76), respectively, in appendix A. As can be seen, the previous approximations and the $[0/1]_{w_{\mathrm{DE}}}(T)$ approximant based approximations have similar relative errors for the physical observables $\Omega_{\mathrm{m}}$ and $E_{\phi}$, with respective $\sim 3\%$ and $\sim 1.5\%$ maxima, while the $[1/1]_{w_{\mathrm{DE}}}(T)$ approximant yields a maximum relative error of $\sim 0.1\%$ for $\Omega_{\mathrm{m}}$ and an even smaller error for $E_{\phi}$, making this approximation observationally indistinguishable from numerical results. Note that changing $\phi_{*}$ so that $\lambda_{*}$ decreases also decreases the errors, which is due to slower rolling of the scalar field. Note also that the errors of all approximations decrease as $z$ increases, which should not come as a surprise since decreasing $z$ entails increased matter domination. Investigation of several other scalar field potentials yield similar results.
(a) Relative errors of $w_{\phi}$
(b) Relative errors of $\Omega_{\mathrm{m}}$
(c) Relative errors of $E_{\phi}$
Figure 1: Thawing quintessence: Relative errors for $w_{\phi}$, $\Omega_{\mathrm{m}}$, $E_{\phi}$. The dotted curves depict the relative errors of the new $[0/1]_{w_{\mathrm{DE}}}(T)$ based approximations (59) while the thick dotted curves represent the relative errors of the $[1/1]_{w_{\mathrm{DE}}}(T)$ based approximations (60). In Figure (a) the dashed curve is the approximation by Scherrer and Sen [52] given in eq. (70), while the space-dashed curve is the approximation obtained by Chiba [53], see eq. (76). In Fig. (b) and (c) the dashed curves correspond to the $\Lambda$CDM solution given in (19), which is the basis for the approximations of both Scherrer and Sen [52] and Chiba [53].
Tracking quintessence
As our example for tracking quintessence we choose the inverse power law potential, for which
$V=V_{0}\phi^{-p},\qquad\lambda=\frac{p}{\phi},\qquad\Gamma=1+\frac{1}{p}=\mathrm{constant}.$ (64)
Thus the derivatives of $\Gamma$ are zero and hence $\Gamma^{(1)}=0$ and $\Gamma^{(2)}=0$ in (57), while, as for all the present tracking quintessence models, $w_{\infty}=-2/(2+p)$. Thus, $w_{\infty}$, $\gamma$ and $\beta$ only involve a single independent parameter, $p$, or, equivalently, $w_{\infty}$, which yields a slightly simpler expressions for $\gamma$ and $\beta$ than $p$. Using the system (52) and the value $\Omega_{\phi 0}=0.68$, which form a surface in the $(\tilde{\phi},u,v)$ state space that intersects the tracker orbit at a point found numerically, and where the numerical calculations also give us $w_{\phi}=u^{2}-1$, $\Omega_{\mathrm{m}}=1-\Omega_{\phi}=1-3\tilde{\phi}v^{2}$, while $E_{\phi}$ is found numerically by integrating [^16] The present model, like all scalar field models, also admits a 1-parameter set of thawing quintessence solutions in the state space, see [32], which we for brevity do not discuss.
$E_{\phi}^{\prime}=-\frac{3}{2}\left(1+w_{\mathrm{tot}}\right)E_{\phi}=-\frac{3}{2}\left(1+w_{\phi}\Omega_{\phi}\right)E_{\phi}=-\frac{3}{2}\left[1+3\tilde{\phi}v^{2}(u^{2}-1)\right]E_{\phi}.$ (65)
[…]
(a) Relative errors of $w_{\phi}$
(b) Relative errors of $\Omega_{\mathrm{m}}$
(c) Relative errors of $E_{\phi}$
Figure 2: Tracking quintessence: Relative errors for $w_{\phi}$, $\Omega_{\mathrm{m}}$, $E_{\phi}$. The dotted curves depict the relative errors for the $[0/1]_{w_{\mathrm{DE}}}(T)$ based approximations given by (59), while the thick dotted curves represent the relative errors for the $[1/1]_{w_{\mathrm{DE}}}(T)$ based approximations (60). In Figure (a) the space-dashed and dashed curves describe the Chiba approximation [54], given in (79), truncated at 1st and 2nd order in $\Omega_{\phi}$, respectively. In Fig. (b) and (c) the dashed curves depict the $w$CDM solution given in (19), which form the basis for the Chiba approximations.
5 Discussion
In this paper we have introduced simple and accurate approximations for thawing and tracking quintessence situated in a broader DE context. This was accomplished by identifying that thawing and tracking quintessence are associated with ‘attractor solutions’ for which $\lim_{N\rightarrow-\infty}\Omega_{\phi}=0$ (and hence $\lim_{N\rightarrow-\infty}\Omega_{\mathrm{m}}=1$) and $\lim_{N\rightarrow-\infty}w_{\phi}=w_{\infty}$, where $w_{\infty}=-1$ for thawing quintessence while $-1<w_{\infty}<0$ for tracking quintessence. These asymptotic properties served as motivation for situating the results in a more general DE setting, which resulted in series expansions in $T=T_{0}\exp(-3w_{\infty}N)$. To improve the accuracy we used Padé approximants where the $[1/1]$ Padé approximant for $\Omega_{\mathrm{DE}}$ was used to relate the approximations to the present time initial value $\Omega_{\mathrm{DE}0}$, or, equivalently, $\Omega_{\mathrm{m}0}=\rho_{\mathrm{m}0}/3H_{0}^{2}$.
We note that the $[0/1]_{w_{\mathrm{DE}}}(T)$ Padé approximant contains as many model parameters ($w_{\infty}$ and $\gamma$) as CPL ($w_{0}$, or equivalently $w_{\infty}$, and $w_{a}$) while the $[1/1]_{w_{\mathrm{DE}}}(T)$ approximant has an additional model parameter $\beta$. However, some potentials yield parameter relationships (or one can impose relationships in the DE context) between $w_{\infty}$, $\gamma$ and $\beta$; notably the inverse power law potential, $V\propto\phi^{-p}$, yields $w_{\infty}(p)$, $\gamma(p)$, $\beta(p)$ and thus our approximations for these models contain one less model parameter than CPL; note, however, that to obtain a specific spacetime, one, in addition to model parameters, needs to specify the initial data $H_{0}$ and $\rho_{\mathrm{m}0}$, or some equivalent initial data, e.g. $H_{0}$ and $\Omega_{\mathrm{DE}0}$.
In this paper we have followed the majority of work in this area by focussing on generating results from $w_{\mathrm{DE}}$, but there are many other possibilities, as well as further developments and applications, where we describe some in the following list:
  • Instead of using $w_{\mathrm{DE}}$ as the starting point, one can use other variables. For example, as pointed out in footnote 7, our series expansions also give rise to a series expansion in $T$ for $E_{\mathrm{DE}}^{2}$ and thereby also for $E_{\mathrm{DE}}$. This expansion can subsequently be used to obtain various Padé approximants, which can then serve as the starting point for calculating other quantities. Or one can derive series expansions for one of the distance measures and use this for constructing Padé approximants, from which other variables and observables can be derived.
  • We have here neglected radiation, which, however, can be included in several different ways; one can e.g. use the fixed points on the radiation boundary as the starting point instead of the fixed points on the matter boundary when it comes to quintessence, as was done in [40]. Or one can simply add radiation, $\rho_{\gamma}=\rho_{\gamma 0}\exp(-4N)=3H_{0}^{2}\Omega_{\gamma 0}\exp(-4N)$, to $\rho$, i.e. $\rho=\rho_{\gamma}+\rho_{\mathrm{m}}+\rho_{\mathrm{DE}}$, and keep the present $w_{\mathrm{DE}}$ expressions, since radiation has a negligible influence on quintessence evolution, and, e.g., use that $3H^{2}=\rho$.
  • For some purposes there might be of interest to obtain an expression for how the scalar field changes during thawing quintessence from the past freezing value $\phi_{*}$ to some arbitrary $e$-fold time $N$. This can be done in several ways, but constructing the $[1,1]_{\phi^{\prime}}$ Padé approximant on the righ hand side of $\phi^{\prime}$ in (43b), i.e. $\sqrt{3(\gamma-1)}\,T/(1+\sigma T)$ where $\sigma=\frac{4}{5}\left(1+\frac{\eta_{*}}{6}\right)$, and then integrating from $\infty$- ∞ to $N$ yields an excellent approximation [^17] It is, e.g., a better approximation than that of Raveri et al. [55] given in equation (72) in appendix A. given by
    $\phi=\phi_{*}+\sqrt{(\gamma-1)/3}\left(\frac{1}{\sigma}\right)\ln{\left(1+\frac{\sigma\Omega_{\phi 0}}{(1-\gamma\Omega_{\phi 0})e^{-3N}}\right)}.$ (66)
  • One can use other dynamical systems variables than the present ones to obtain somewhat different results. For example, we could have used the variables in [32] to deal with tracking quintessence, which would have been useful for the hyperbolic potentials that in addition to the inverse power law potential was dealt with in that paper. [^18] This would had come with the price that $\Gamma$ would not have appeared explicitly in the results for $\gamma$ and $\beta$ due to the different regularization of $\lambda$. However, to improve convergence one should then replace the scalar field variable $\bar{\varphi}$ in [32] with $\tilde{\varphi}=\bar{\varphi}^{2}$.
  • In the spirit of CPL, it is possible to improve our quintessence approximations even further by viewing $\gamma$ and $\beta$ as free parameters, not determined by the scalar field potential. This, however, means that one only would do curve fitting, which implies less theoretical content and predictive power.
  • We have here focussed on obtaining thawing and tracking quintessence approximations situated in a broader DE context for spatially homogeneous, isotropic and flat spacetimes. A natural next step is to use these approximations in order to interpret and assess the rapidly growing set of increasingly precise observational data. In particular, the present results, or variations thereof, can serve as input in cosmological perturbation theory to bring the results closer to a wide set of observations.
  • We finally note that the present approach and variations thereof can easily be applied to other models and theories, especially for those admitting an Einstein frame, although this usually adds parameters that diminish predictive power.
Acknowledgments
A. A. is supported by FCT/Portugal through CAMGSD, IST-ID, projects UIDB/04459/2020 and UIDP/04459/2020, and by the H2020-MSCA-2022-SE project EinsteinWaves, GA No. 101131233. A.A. would also like to thank the CMA-UBI in Covilhã for kind hospitality. C. U. would like to thank the CAMGSD, Instituto Superior Técnico in Lisbon for kind hospitality.
Appendix A Quintessence approximations and comparisons
In this appendix we review the most prominent quintessence approximations in the literature, including some simplified derivations and extensions; moreover, we will present them in a more unified and concise manner than their original form. We begin with thawing quintessence.
A.1 Thawing quintessence
Let us first derive an approximation introduced by Sherrer and Sen (2008) in [52] and further explored by e.g. Agrawal et al. [56] and Raveri et al. [55]. Instead of using the state vector $(\phi,\Sigma_{\phi},\Omega_{\phi})$ we construct a dynamical system closely connected with $\tilde{\phi},u,v$ based on the state vector $(\phi,u,\omega)$ where [^19] The variable $u$ was introduced in [30] while the variable $v$ in that paper is just $v=\omega/\sqrt{3}$.
$u=\Sigma_{\phi}\sqrt{\frac{2}{\Omega_{\phi}}}=\frac{\phi^{\prime}}{\sqrt{3\Omega_{\phi}}},\qquad\omega=\sqrt{\Omega_{\phi}},$ (67)
which due to (39) yields the dynamical system
${\phi}^{\prime}$ $=\sqrt{3}u\omega,$ (68a)
$u^{\prime}$ $=\frac{1}{2}(2-u^{2})[\sqrt{3}\lambda({\phi})\omega-3u],$ (68b)
$\omega^{\prime}$ $=\frac{3}{2}(1-u^{2})\omega(1-\omega^{2}).$ (68c)
Next $\phi$ is assumed to take a frozen constant value $\phi=\phi_{*}$ while the two last equations in (68) are linearized in $u$ only, which results in the following equations for $u$ and $\omega$:
$u^{\prime}$ $\approx-3u+\sqrt{3}\lambda_{*}\omega,$ (69a)
$\omega^{\prime}$ $\approx\frac{3}{2}\omega(1-\omega^{2}),$ (69b)
where $\lambda_{*}=\lambda({\phi}_{*})$. The second equation, which corresponds to $\Omega_{\phi}^{\prime}\approx 3\Omega_{\phi}(1-\Omega_{\phi})$ and hence $[\ln\{\Omega_{\phi}/(1-\Omega_{\phi})\}]^{\prime}=[\ln(\Omega_{\phi}/\Omega_{\mathrm{m}})]^{\prime}=[\ln(\rho_{\phi}/\rho_{\mathrm{m}})]^{\prime}\approx 3$, where $\rho_{\mathrm{m}}=\rho_{\mathrm{m},0}\exp(-3N)$ yields $\rho_{\phi}\propto\mathrm{constant}$ and thereby the $\Lambda$CDM expression for $\omega(N)$. Inserting this solution into (69a) and integrating gives $u(N)$, from which $w_{\phi}=u^{2}-1$ follows; deparameterizing the solution results in $u(\omega)$ and thereby $w_{\phi}(\omega)$; alternatively, solve directly $du/d\omega=u^{\prime}/\omega^{\prime}\approx-\frac{2u}{\omega(1-\omega^{2})}+\frac{2\lambda_{*}}{\sqrt{3}(1-\omega^{2})}$:
$\begin{split}1+w_{\phi}&=\frac{\lambda_{*}^{2}}{3}\left[\omega^{-1}-(\omega^{-2}-1)\tanh^{-1}(\omega)\right]^{2}\\ &=\frac{\lambda_{*}^{2}}{3}\left[\Omega_{\phi}^{-1/2}-(\Omega_{\phi}^{-1}-1)\tanh^{-1}(\Omega_{\phi}^{1/2})\right]^{2},\end{split}$ (70a)
$\Omega_{\phi}$ $=\left[1+\left(\Omega_{\phi 0}^{-1}-1\right)e^{-3N}\right]^{-1},$ (70b)
where the last expression coincides with that of $\Lambda$CDM when identifying $\Omega_{\phi}$ with $\Omega_{\Lambda}$ (this approximation thereby does not give thawing corrections to the line element since it is just the $\Lambda$CDM line element); the above expressions summarize eqs. (23), (25) and (26) in Scherrer and Sen [52]. Taken together with eq. (68a) and $u^{2}=1+w_{\phi}$ this results in
$\phi^{\prime}=\lambda_{*}\left[1-(\omega^{-1}-\omega)\tanh^{-1}(\omega)\right]=\lambda_{*}\left[1-(\Omega_{\phi}^{-1/2}-\Omega_{\phi}^{1/2})\tanh^{-1}(\Omega_{\phi}^{1/2})\right],$ (71)
where inserting (70b) yields $\phi^{\prime}(N)$. This expression agrees with eq. (15) in Agrawal et al. [56] for the exponential potential where $\lambda$ is a constant. Integration, with the condition that ${\phi}={\phi}_{*}$ when $N\rightarrow-\infty$ and $\Omega_{\varphi}\rightarrow 0$ (using $\phi^{\prime}=\Omega_{\phi}^{\prime}d\phi/d\Omega_{\phi}$ and $\Omega_{\phi}^{\prime}\approx 3\Omega_{\phi}(1-\Omega_{\phi})$), results in
$\phi\approx\phi_{*}+\frac{2\lambda_{*}}{3}\left[\Omega_{\phi}^{-1/2}\tanh^{-1}(\Omega_{\phi}^{1/2})-1\right],$ (72)
where $\phi(N)$ is obtained by inserting (70b). This approximation coincides with eq. (8) in Raveri et al. [55].
Other types of thawing quintessence approximations in the literature are based on series expansions of $V(\phi)$ at some $\phi_{*}$, and thus also of $\lambda(\phi)$ when the latter is regular. An example of this were given by Cahn et al. (2008) [57] who gave approximate expressions for $\phi^{\prime}$ (replacing $\dot{\phi}$ with $\phi^{\prime}$ in their eq. (15)) and $\phi$, although note that $\phi_{*}$ is missing in their eq. (16), thus hiding that you typically are interested in a domain of $\phi_{*}$. These expressions correspond to the present second order expressions in eqs. (43a) and (43b), whose accuracy we subsequently improved by taking Padé approximants.
Based on earlier results by Dutta and Scherrer (2008) [58], Chiba (2009) [53] Taylor expanded the potential to second order in $\phi-\phi_{*}$. By a sequence of ad hoc approximations in the proper time $t$ the author derived a series of results (see p. 083517-4 in [53]), which can be extended to also include $\phi^{\prime}$ and $w_{\mathrm{tot}}$, more simply expressed in $\omega=\sqrt{\Omega_{\phi}}$ as follows:
$\phi$ $=\phi_{*}+\frac{2\lambda_{*}}{3}\left(\frac{(1-\omega^{2})^{\frac{1}{2}(1-K)}\left[(1+\omega)^{K}-(1-\omega)^{K}\right]-2K\omega}{K\omega(K^{2}-1)}\right)$ (73a)
$\phi^{\prime}$ $=\lambda_{*}\frac{(1-\omega^{2})^{\frac{1}{2}(1-K)}\left[(1+K\omega)(1-\omega)^{K}-(1-K\omega)(1+\omega)^{K}\right]}{K\omega(K^{2}-1)}$ (73b)
$1+w_{\phi}$ $=\frac{\lambda^{2}_{*}}{3}(1-\omega^{2})^{1-K}\left[\frac{(1+K\omega)(1-\omega)^{K}-(1-K\omega)(1+\omega)^{K}}{K\omega^{2}(K^{2}-1)}\right]^{2}$ (73c)
$w_{\mathrm{tot}}$ $=w_{\phi}\omega^{2},$ (73d)
where
$K=\sqrt{1-\frac{4}{3}\eta_{*},}$ (74)
while $\omega=\sqrt{\Omega_{\phi}}$ is given by its $\Lambda$CDM expression
$\omega=\left(1+(\Omega_{\phi 0}^{-1}-1)e^{-3N}\right)^{-1/2}.$ (75)
The right hand side for $\phi^{\prime}$ was obtained by using $\phi^{\prime}=\frac{d\phi}{d\omega}\omega^{\prime}$ and the $\Lambda$CDM relation $\omega^{\prime}=\frac{3}{2}(1-\omega^{2})\omega$. Taking the limit $K\rightarrow 1$ of equations (73a) and (73b), which corresponds to neglecting the second order Taylor expansion term of the potential, i.e. setting $V_{,\phi\phi}(\phi_{*})=0$, leads to an indeterminacy that can be solved by L’Hôpital’s rule. In [53] the author also showed how to convert the expression for $w_{\phi}$ to the case when $K^{2}<0$, which resulted in the approximation
$1+w_{\phi}=\frac{4}{3}\lambda^{2}_{*}(1-\omega^{2})\left[\frac{\tilde{K}\omega\cos{\left(\tilde{K}\sinh^{-1}{\left(\frac{\omega}{\sqrt{1-\omega^{2}}}\right)}\right)}-\sin{\left(\tilde{K}\sinh^{-1}{\left(\frac{\omega}{\sqrt{1-\omega^{2}}}\right)}\right)}}{\tilde{K}\omega^{2}(1+\tilde{K}^{2})}\right]^{2},$ (76)
where
$\tilde{K}=\sqrt{\frac{4}{3}\eta_{*}-1}.$ (77)
In addition, Chiba [53] discussed the approximation given by Crittenden et al. (2007) [59]. Examining the approximations for a variety of potentials we do not find that $K(\phi_{*})\neq 1$ gives better approximations than the simpler $K=1$ expressions, which presumably is due to error cancellations and accumulations in a rather uncontrolled manner during the series of ad hoc approximations that resulted in the $K\neq 1$ expressions.
A.2 Tracking quintessence
Here we consider the tracking quintessence approximations given by Chiba (2010) [54], which in turn relied on earlier work by Watson and Scherrer (2003) [60], for the $\Gamma=\mathrm{constant}$ case, and, in particular, $\Gamma=\mathrm{constant}>1$, which yields the inverse power law potential. This work is based on a linearized equation for a perturbation of $w_{\phi}$ with respect to its asymptotic value $w_{\infty}$. To obtain an improved approximation for the linearized solution for $w_{\phi}$ Chiba [54] considered a series expansion in $e^{-3wN}$, when expressed in $e$-fold time and the present notation. This is the same series expansion quantity as in the present paper, but in contrast to Chiba [54] we used it to produce approximations for the full non-linear equations; thus, the expansions in the present work coincide with those in [60] and [54] for the quantities that have been computed in those two references up to first order but not for higher orders since Chiba [54] only deals with the linearized equation for $w_{\phi}$. Moreover, there is translation freedom in $N$. In the present work this freedom was fixed by the $[1/1]$ Padé approximant for $\Omega_{\phi}$, which resulted in an expansion in $T=\Omega_{\phi 0}e^{-3w_{\infty}N}/(1-\gamma\Omega_{\phi 0})$; Chiba [54] on the other hand fixed this freedom by, in effect, setting $w_{\phi}=w_{\infty}$, to obtain the following approximation for $\Omega_{\phi}$:
$\Omega_{\phi}\approx\frac{\tilde{T}}{1+\tilde{T}}=\frac{\Omega_{\phi 0}}{\Omega_{\phi 0}+(1-\Omega_{\phi 0})e^{3w_{\infty}N}},\qquad\tilde{T}=\left(\frac{\Omega_{\phi 0}}{1-\Omega_{\phi 0}}\right)e^{-3w_{\infty}N},$ (78)
which in turn yields an approximation for $E_{\phi}$, while the series expansion in $\tilde{T}$ for the linearized equation for $w_{\phi}$ resulted in
$w_{\phi}=w_{\infty}+w_{\infty}(1-w_{\infty}^{2})\sum_{n=1}^{\infty}\left(\frac{(-1)^{n-1}}{2n(n+1)w_{\infty}^{2}-(n+1)w_{\infty}+1}\right)\tilde{T}^{n}.$
Note that the quantity $\tilde{T}$ is related to $T$ according to
$\tilde{T}=\left(\frac{1-\gamma\Omega_{\phi 0}}{1-\Omega_{\phi 0}}\right)T,$
which corresponds to a translation in $N$ with the constant $-(1/3w_{\infty})\ln[(1-\gamma\Omega_{\phi 0})/(1-\Omega_{\phi 0})]$.
As a next step Chiba [54] replaced $\tilde{T}$ with $\tilde{T}=\Omega_{\phi}/(1-\Omega_{\phi})$, and expanded the resulting expression for $w_{\phi}$ in $\Omega_{\phi}$, which yields the following series, here truncated at second order, in $\Omega_{\phi}$:
$w_{\phi}\approx w_{\infty}+\left(\frac{w_{\infty}(1-w_{\infty}^{2})}{4w_{\infty}^{2}-2w_{\infty}+1}\right)\Omega_{\phi}\left[1+\left(\frac{w_{\infty}(8w_{\infty}-1)}{12w_{\infty}^{2}-3w_{\infty}+1}\right)\Omega_{\phi}\right].$ (79)
Note that this is an expansion that is designed to improve the solution for the linear equation for $w_{\phi}$, not the actual solution of the non-linear equation for $w_{\phi}$. The improved linear approximation only improves the actual solution to the extent the linear approximation describes that solution, which, fortunately, to some extent is the case for the inverse power law potential, see e.g., Figure 1 in [28]. Not surprisingly, these approximations yield poorer accuracy than the present ones that are based on approximating the full non-linear equations.
Appendix B Scaling freezing quintessence
In [30] it was demonstrated that in appropriate variables, resulting in a global 3D state space setting, scaling freezing quintessence is associated with a single attractor orbit that constitutes the unstable manifold of an isolated fixed point. The situation and key attractor mechanism is thereby the same as for tracking quintessence. An open set of nearby orbits to the (hyperbolic) saddle fixed point is pushed toward the attractor orbit by a 2D invariant unstable manifold boundary set, for details, see [30] and [32]. The difference with tracking quintessence is that scaling freezing quintessence arises from potentials that have an asymptotic exponential behaviour (that is sufficiently steep).
We therefore consider the situation where $\lim_{\phi\rightarrow-\infty}\lambda=\lambda_{-}$, $\lambda_{-}>\sqrt{3}$, but where compatibility with nuclear synthesis data requires $\lambda_{-}\gg\sqrt{3}$, see [30] and references therein. Let us further assume that the scalar field, during the time epoch $N\in(-\infty,0]$ for the attractor solution, rolls down a scalar field potential with $\lambda(\phi)>0$ where $\lambda(\phi)$ is monotonically decreasing, i.e., we assume that
$\Gamma=\frac{V\,V_{,\phi\phi}}{V_{,\phi}^{2}}=1+(\lambda^{-1})_{,\phi}>1\quad\text{and}\quad\lim_{\phi\rightarrow-\infty}\Gamma=1.$ (80)
For our analysis we will consider the dynamical system for the state space vector $(\lambda,\Sigma_{\phi},\Omega_{\phi})$, which can be obtained from (39):
$\lambda^{\prime}$ $=-\sqrt{6}(\Gamma-1)\lambda^{2}\Sigma_{\phi},$ (81a)
$\Sigma_{\phi}^{\prime}$ $=-\frac{3}{2}(1+\Omega_{\phi}-2\Sigma_{\phi}^{2})\Sigma_{\phi}+\sqrt{\frac{3}{2}}\,\lambda\,(\Omega_{\phi}-\Sigma_{\phi}^{2}),$ (81b)
$\Omega^{\prime}_{\phi}$ $=3(\Omega_{\phi}-2\Sigma_{\phi}^{2})(1-\Omega_{\phi}),$ (81c)
where $\Gamma$ now is considered to be a function of $\lambda$, i.e., $\Gamma=\Gamma(\lambda)$.
Before continuing, let us note that the simple example of a potential consisting of two exponential terms,
$V=V_{-}e^{-\lambda_{-}\phi}+V_{+}e^{-\lambda_{+}\phi},\qquad\lambda_{-}>\lambda_{+},$ (82)
yields
$\Gamma=1+\frac{(\lambda-\lambda_{-})(\lambda_{+}-\lambda)^{2}}{\lambda^{2}}.$ (83)
In the present variables, scaling freezing quintessence is associated with the unstable manifold orbit that originates from the fixed point
$\mathrm{S}\!:\quad(\lambda,\Sigma_{\phi},\Omega_{\phi})=\left(\lambda_{-},\sqrt{\frac{3}{2}}\left(\frac{1}{\lambda_{-}}\right),\frac{3}{\lambda^{2}_{-}}\right).$ (84)
Next we assume that $\Gamma(\lambda)$ can be Taylor expanded around $\lambda_{-}$, i.e.,
$\Gamma=1+\Gamma_{1}(\lambda-\lambda_{-})+\frac{\Gamma_{2}}{2}(\lambda-\lambda_{-})^{2}+\dots,\qquad\Gamma_{n}=\left.\frac{d^{n}\Gamma}{d\lambda^{n}}\right|_{\lambda=\lambda_{-}}.$ (85)
Note that (83) is regular at $\lambda=\lambda_{-}$ for the potential with two exponential terms and that
$\Gamma_{1}=\left(1-\frac{\lambda_{+}}{\lambda_{-}}\right)^{2},\qquad\Gamma_{2}=\frac{2}{\lambda_{-}}\left(1-\left(\frac{\lambda_{+}}{\lambda_{-}}\right)^{2}\right)$ (86)
in this case
We can now proceed in the same manner as for thawing and tracking quintessence and first make an expansion based on the unstable eigenvalue, $-3\lambda_{-}\Gamma_{1}$, which leads to expansions in
$T=Ce^{-3\lambda_{-}\Gamma_{1}N}.$ (87)
The fact that $\lim_{N\rightarrow-\infty}\Omega_{\phi}>0$ and that the eigenvalue involves two parameters, $\lambda_{-}$ and $\Gamma_{1}$, result in quite messy expressions for the series expansions. For brevity we will therefore truncate the series already at the linear order in $T$:
$\lambda$ $=\lambda_{-}-T+\dots,$ (88a)
$\Sigma_{\phi}$ $=\sqrt{\frac{3}{2}}\left(\frac{1}{\lambda_{-}}\right)\left(1+\alpha T+\dots\right),$ (88b)
$\Omega_{\phi}$ $=\frac{3}{\lambda^{2}_{-}}\left(1+\beta T+\dots\right),$ (88c)
where
$\alpha=\frac{\lambda^{3}_{-}\Gamma_{1}+\lambda^{2}_{-}-3}{\lambda_{-}(2\lambda^{4}_{-}\Gamma^{2}_{1}-\lambda^{3}_{-}\Gamma_{1}+\lambda^{2}_{-}-3)},\qquad\beta=\frac{2(\lambda^{2}_{-}-3)}{\lambda_{-}(2\lambda^{4}_{-}\Gamma^{2}_{1}-\lambda^{3}_{-}\Gamma_{1}+\lambda^{2}_{-}-3)}.$ (89)
Using that $w_{\phi}=-1+2\Sigma_{\phi}^{2}/\Omega_{\phi}$ yields
$w_{\phi}=(2\alpha-\beta)T+\dots,$ (90)
where
$2\alpha-\beta=\frac{2\lambda^{2}_{-}\Gamma_{1}}{\lambda_{-}(2\lambda^{4}_{-}\Gamma^{2}_{1}-\lambda^{3}_{-}\Gamma_{1}+\lambda^{2}_{-}-3)}.$ (91)
Note that $\lim_{N\rightarrow-\infty}w_{\phi}=0$, which leads to that asymptotically toward the past $\rho_{\phi}\propto\rho_{\mathrm{m}}\propto a^{-3}$, while $w_{\phi}$ decreases toward the future, hence the name scaling freezing.
We leave the higher order series expansions, the construction of Padé approximants, the identification of $C$ with present day data, and subsequent investigations, such as error estimates for various potentials, for the interested reader.